The average ago of the starting line-up on an NBA basketball team is 25 years. If four of the five players are 21,23,24, and 31 years old, find the age of the fifth player.
Answer by checkley77(12844)

21,23,24, and 31 years old,
(21+23+24+31+x)/5=25
(99+x)/5=25
99+x=5*25
x=125-99
x=26 years is the age of the fifth player.
Proof:
(21+23+24+31+26)/5=25
125/5=25
125=5*25
125=125